OpenStudy (anonymous):

which of the following equations describes the line shown below? Check all that apply.

OpenStudy (anonymous):

A. y - 4 = 3 (x - 2) B. y - 13 = 3(x - 1) C. y + 2 = 3(x - 4) D. y - 2 = 3(x - 4) E. y - 4 = 3(x + 2) F. y - 1 = 3(x - 13)

OpenStudy (anonymous):